OpenStudy (anonymous):

ok so i need help finding the perimeter, the hight,the surface aria, and the slant hight of a square pyramid..who can help

OpenStudy (anonymous):

Pythagoras probably could have helped :(

OpenStudy (anonymous):

what

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pythagorean_theorem

OpenStudy (anonymous):

well i got that...i just have a limited about of numbers that help me at all

OpenStudy (anonymous):

i have a volume that =340 and values of x1-12 that equal the base lengths

OpenStudy (anonymous):

Is it a regular square based pyramid (i.e. the 4 triangles are equilateral)? If so, then that is probably enough. If not, you're f*cked

OpenStudy (anonymous):

umm...thanks

OpenStudy (anonymous):

I was trying to help. No need to get all dorky on me.

OpenStudy (anonymous):

lol...i am a dork to begin with...so im sorry but you are going to have to retard this out for me

OpenStudy (anonymous):

Have fun.

OpenStudy (anonymous):

not helping..im having issues trying to figure out what goes in the calculator

OpenStudy (anonymous):

What is x1? just a random number?

OpenStudy (anonymous):

ok so i got the perimeter which is \[p=x \times 4 \]

OpenStudy (anonymous):

now i need to get the side length which is \[a^2+b^2=c^2 so what is a b and c\]

OpenStudy (anonymous):

OK I will try and do this quickly - what EXACTLY do you know?

OpenStudy (anonymous):

i know the volume is 340 inches cubed..and that 1 is the side length

OpenStudy (anonymous):

Do you mean 1 is the length of the 4 edges of the base? Or the length of the slants?

OpenStudy (anonymous):

oops yes that is what i ment

OpenStudy (anonymous):

Well if that is the case, the volume = 340 = (1/3) * 1^2 * height , so height = 1020... which sounds retardedly big. But if that was right, the slant length ^2 = 1^2 + 1020^2 And surface area = 2 * slant length * base length + base length squared Weird question, but I have to bye. bye babe xx